Suppose Standure is a monopoly firm in the market for 3D online games in UK. The marginal cost functions for this firm is giv(b) Determine the producer surplus Standure has under the conditions described above. (4 marks)(c) What is the value of deadweight loss if Standure acts as a monopolist? (10 marks)
Answer #1

(@) under perfect competition, equilibrium occurs where Р: Mc = 150-Q = 0.25Q + 15 → 625Q = 135 Q* - 108 Ans p*= 150-108 = 42